OpenStudy (anonymous):

Evaluate the expression for x = 2, y = 1, and z = -3. Simplify. 2x2+y+10z I got -21 am I right?

OpenStudy (anonymous):

2(2) ^ 2 + 1 + 10(-3) \[2(2)^2 + 1 + 10(-3) = 8+1-30\] =-21 Yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thank you!